OpenStudy (anonymous):

Ten inches of heavy, wet snow are equivalent to one inch in rain. Estimate the water content in 14 inches of heavy, wet snow.

OpenStudy (luigi0210):

.1x=y x=inches of heavy wet snow y=rain in inches

OpenStudy (anonymous):

I have 10 = 14 --- --- 1 x 10x= 1*14 10x= 14 10x= 14 x= 1.4

OpenStudy (texaschic101):

proportion time 10/1 = 14/x (10 in to 1 in = 14 in to x in) cross multiply 10x = 14 x = 14/10 x = 7/5 = 1.4 ANSWER : 1.4 inches

OpenStudy (anonymous):

Thank you

OpenStudy (luigi0210):

And you were right again

OpenStudy (texaschic101):

thats what we are here for :)